# CausalCite: A Causal Formulation of Paper Citations

### v2

@@ 1 Introduction
A traditional approach to recognize paper quality is peer review, a mechanism that requires large efforts, and yet has inherent randomness and flaws (Cortes and Lawrence, 2021; Rogers et al., 2023; Shah, 2022; Prechelt et al., 2018; Resnik et al., 2008). Moreover, the number of papers after peer review is still overwhelmingly large for researchers to read, leaving the challenge of identifying truly impactful research unaddressed. Another commonly used metric is citations. However, this metric faces criticism for biases, such as a preference for survey, toolkit, and dataset papers (Zhu et al., 2015; Valenzuela-Escarcega et al., 2015). Citations, together with altmetrics (Wilsdon et al., 2015), which incorporates social media attention to a paper, often have biases towards papers with extensive publicity and promotion, or those authored by established figures in the field.
@@ 3.2 Can Existing Causal Inference Methods Handle This Control?
$\widehat{\mathrm{ACI}}(a)=\sum_{\bm{x}}P(\bm{x})\left(\frac{1}{|\mathcal{T}_{\bm{x}}|}\sum_{i\in\mathcal{T}_{\bm{x}}}y_{i}-\frac{1}{|\mathcal{C}_{\bm{x}}|}\sum_{j\in\mathcal{C}_{\bm{x}}}y_{j}\right)~{},$
@@
To handle adjustment variables in the high-dimensional space, one possible way is to avoid pre-defining all their possible intervals, but, instead, iterating over each unit in the treatment group to match for its closest control unit. Consider a given follow-up paper $b$, and a set of candidate control papers $\bm{C}$, where each paper $c_{i}$ has a citation count $y_{i}$, and vector representation $\bm{t}_{i}$ of the confounders (e.g., research topic). One-to-one matching estimates PCI as
$\begin{split}\widehat{\mathrm{PCI}}(a,b)&=y_{b}-y_{\operatorname*{argmax}_{c_{i}\in\bm{C}}m_{i}}\\ &=y_{b}-y_{\operatorname*{argmax}_{c_{i}\in\bm{C}}\mathrm{sim}(\bm{t}_{y},\bm{t}_{i})}~{},\end{split}$ (4)
@@
3.3 How Do We Extend Causal Inference to Text Variables?
@@ 3.3.1 Theoretical Formulation of TextMatch: Stabilizing Text Matching by Synthesis
$\widehat{\mathrm{PCI}}(a,b)=y_{b}-\sum_{c_{i}\in\bm{C}}w_{i}y_{i}=y_{b}-\sum_{c_{i}\in\bm{C}}\frac{m_{i}}{\sum_{i}m_{i}}y_{i}~{},$ (5)
@@ 3.3.3 Key Challenges and Mitigation Methods
There are several technical challenges, which we address below.
@@
Some Mediators Might Be Entangled in the Confounder Variable Another key challenge to highlight is that the text variable we use for the confounder might accidentally include some mediator information. For example, the quality or performance of a paper could be expressed in the abstract, such as “we achieved 90% accuracy.” Therefore, we conduct a specific preprocessing procedure before feeding the text variable to the similarity function. For the $\mathrm{RemoveMediator}$ function in 19, we exclude all numerical expressions such as percentage numbers, as well as descriptions such as “state-of-the-art.” For generalizability, the essence of this step is a entanglement action to separate the confounder variable (in this case, the research content) and all the descendants of the treatment variable (in this case, mentions of the performance). For more complicated cases in future work, we recommend a separate disentanglement model to be applied here.
@@
Overcoming Numerical Estimation Challenges Given the large number of papers, it is numerically challenging to aggregate the TCI from individual PCIs, because the number of follow-up papers for a study can be up to tens of thousands, such as the 57,200 citations by 2023 for the ImageNet paper (Deng et al., 2009). To avoid extensively running PCI for all follow-up papers, we propose a new numerical estimation method using a carefully designed random paper subset.
@@ 4.2 Author-Identified Paper Impact
Figure 4 reports the accuracy of our CausalCite metric, together with two existing citation metrics: citations, and SSHI citations (Valenzuela-Escarcega et al., 2015). See the detailed derivation of the accuracy scores in Section C.2. From this table, we can see that our CausalCite metric achieves the highest accuracy, 80.29%, which is 5 points higher than SSHI, and 9 points higher than the traditional citations.
@@ 4.3 Test-of-Time Paper Analysis
In Figure 4, we show the correlations of various metrics with the ToT awards. In this table, CausalCite achieves the highest correlation of 0.639, which is +30.14% better than that of citations. Furthermore, we visualize the correspondence of our metric and ToT, and observe a substantial difference between the CausalCite distributions of ToT vs. non-ToT papers in Figure 4. We also show three examples of ToT papers in Figure 4, where the ToT papers differ from the non-ToT papers by one or two orders of magnitude.
@@ 5.1 Do Best Papers Have High Causal Impact?
5.2 What Is the Curve Shape of the CausalCite Distributions?
@@ 6 Related Work
With the increasing availability of large datasets and the advent of digital technologies, new opportunities for bibliometric analysis have emerged. Iqbal et al. (2021) highlighted the role of natural language processing and machine learning in enhancing in-text citation analysis. Similarly, Umer et al. (2021) explored the use of textual features and SMOTE resampling techniques in scientific paper citation analysis. Jebari et al. (2021) analyzed citation context to detect research topic evolution, showcasing data analysis for scientific discourse. Chang et al. (2023) explored augmenting citations in scientific papers with historical context, offering a novel perspective on citation analysis. Manghi et al. (2021) introduced scientific knowledge graphs, an innovative method for evaluating research impact. Bittmann et al. (2021) explored statistical matching in bibliometrics, discussing its utility and challenges in post-matching analysis. The use of AI in bibliometric analysis is highlighted in research by Chubb et al. (2022) and the systematic review of AI in information systems by Collins et al. (2021). Network analysis approaches, as discussed by Chakraborty et al. (2020) in the context of patent citations and by Dawson et al. (2014) in learning analytics, further illustrate the diverse applications of advanced methodologies in understanding citation patterns.
@@ 7 Conclusion
There are several limitations for our work. For example, as mentioned previously, our metric has a high computational budget. Future work can explore more efficient optimization methods.
@@
Data Collection and Privacy The data used in this work are all from Open Source Semantic Scholar data, with no user privacy concerns. The potential use of this work is for finding papers that are unique and innovative but don’t get enough citations due to loack of popularity or awareness of the field. This metric can act as an aid when deciding impact of papers, but we do not suggest its usage without expert involvement. Through this work, we are not trying to demean or criticize anyone’s work we only intend to find more papers that have made a valuable contribution to the field.
CS Centric Perspective The authors of this paper work in Computer Science (mostly Machine Learning) hence a lot of analysis done on the quality of papers that required sanity checks are done on ML papers. The conferences selected for doing the ToT evaluation were also CS Top conferences, hence they might have induced some biases. The metric in general has been created generically and should be applicable to other domains as well, the Author Identified Most Influential Papers study is also done on a generalized dataset, but we encourage readers in other disciplines to try out the metric on papers from their field.
@@
This project originates as part of the AI Scholar series of projects that Zhijing Jin started since 2021, as she identified that causal inference over papers is a valuable research setting with sufficient data and rich causal phenomenon. Bernhard Schölkopf came up with the formulation that the action of citation itself has a causal nature, and can thus be formulated as a causal inference question. Zhijing, Bernhard, and Siyuan Guo settled down the overall project design.
@@ A.3 Experiment to Select the Best Embedding Method
When selecting the text encoder for our TextMatch method, we compare among the three LLMs pre-trained on scientific papers, SciBERT, MPNet, and SPECTER. Specifically, we conduct a small-scale experiment to see how much the similarities scores based on the embedding of each model align with human annotations. As for the annotation process, we first collect a set of random papers as pivot papers, and for each pivot paper, we identify ten papers, from the most similar to the least, with monotonically decreasing similarity. We collect a total of 100 papers consisting of ten such collections, for which we show an example in Table 6. Then we see how the resulting similarity scores conform to this order by deducting the percentage of papers that are out of place in the ranking.
@@ C.2 Additional Information for the Author-Identified Paper Impact Experiment
- A Quantitative Analysis and Performance Study for Similarity-Search Methods in High-Dimensional Spaces
0 3.519 1777 156
- Optimizing bitmap indices with efficient compression
0 3.519 375 40
- Data Warehouses And Olap: Concepts, Architectures And Solutions
0 3.526 187 11
- Histogram-aware sorting for enhanced word-aligned compression in bitmap indexes
0 3.543 17 1
- CubiST++: Evaluating Ad-Hoc CUBE Queries Using Statistics Trees
0 3.543 5 1
- Improving Performance of Sparse Matrix-Vector Multiplication
0 3.543 114 11
- Binary Gray Codes with Long Bit Runs
0 3.543 53 4
- Analysis of Basic Data Reordering Techniques
0 3.543 16 1
- Tree Based Indexes Versus Bitmap Indexes: A Performance Study
0 3.543 24 0
- Secondary indexing in one dimension: beyond b-trees and bitmap indexes
0 3.543 10 1
- A comparison of five probabilistic view-size estimation techniques in OLAP
0 3.543 24 1
- Compression techniques for fast external sorting
0 3.543 16 0
- A Note on Graph Coloring Extensions and List-Colorings
0 3.543 33 1
- Using Multiset Discrimination to Solve Language Processing Problems Without Hashing
0 3.543 52 2
- Monotone Gray Codes and the Middle Levels Problem
0 3.543 80 5
- The Art in Computer Programming
0 3.543 9242 678
- An Efficient Multi-Component Indexing Embedded Bitmap Compression for Data Reorganization
0 3.543 8 2
- The LitOLAP Project: Data Warehousing with Literature
0 3.543 8 0
- Multi-resolution bitmap indexes for scientific data
0 3.583 96 3
- Notes on design and implementation of compressed bit vectors
0 3.583 81 12
- Compressing Large Boolean Matrices using Reordering Techniques
0 3.595 88 7
- Compressing bitmap indices by data reorganization
1 3.595 53 4
- Model 204 Architecture and Performance
0 3.635 238 10
- On the performance of bitmap indices for high cardinality attributes
1 3.654 196 10
- A performance comparison of bitmap indexes
0 3.655 86 9
- Minimizing I/O Costs of Multi-Dimensional Queries with Bitmap Indices
0 3.692 16 0
- Evaluation Strategies for Bitmap Indices with Binning
0 3.692 69 3
- C-Store: A Column-oriented DBMS
0 3.710 1241 111
- Byte-aligned bitmap compression
0 3.793 209 48
- Bit Transposed Files
0 3.837 84 10
- Space efficient bitmap indexing
0 4.011 96 16

### v3

@@ 1 Introduction
A traditional approach to recognize paper quality is peer review, a mechanism that requires large efforts, and yet has inherent randomness and flaws (Cortes and Lawrence, 2021; Rogers et al., 2023; Shah, 2022; Prechelt et al., 2018; Resnik et al., 2008). Moreover, the number of papers after peer review is still overwhelmingly large for researchers to read, leaving the challenge of identifying truly impactful research unaddressed. Another commonly used metric is citations. However, this metric faces criticism for biases, such as a preference for survey, toolkit, and dataset papers (Zhu et al., 2015; Valenzuela-Escarcega et al., 2015). Together with altmetrics (Wilsdon et al., 2015), which incorporates social media attention to a paper, both metrics also bias towards papers from major publishing countries (Rungta et al., 2022; Gomez et al., 2022), with extensive publicity and promotion, and authored by established figures.
@@ 3.2 Can Existing Causal Inference Methods Handle This Control?
$\widehat{\mathrm{ACI}}(a)=\sum_{\bm{x}}P(\bm{x})\left(\frac{1}{|\mathcal{T}_{\bm{x}}|}\sum_{i\in\mathcal{T}_{\bm{x}}}y_{i}-\frac{1}{|\mathcal{C}_{\bm{x}}|}\sum_{j\in\mathcal{C}_{\bm{x}}}y_{j}\right)~{},$ (4)
@@
To handle high-dimensional adjustment variables, one possible way is to avoid pre-defining all their possible intervals, but, instead, iterating over each unit in the treatment group to match for its closest control unit (e.g., McGue et al., 2010; Sato et al., 2022). Consider a given follow-up paper $b$, and a set of candidate control papers $\bm{C}$, where each paper $c_{i}$ has a citation count $y_{i}$, and vector representation $\bm{t}_{i}$ of the confounders (e.g., research topic). One-to-one matching estimates PCI as
$\begin{split}\widehat{\mathrm{PCI}}(a,b)&=y_{b}-y_{\operatorname*{argmax}_{c_{i}\in\bm{C}}m_{i}}\\ &=y_{b}-y_{\operatorname*{argmax}_{c_{i}\in\bm{C}}\mathrm{sim}(\bm{t}_{b},\bm{t}_{i})}~{},\end{split}$ (5)
@@
3.3 How Do We Extending Causal Inference to Text Variables?
@@ 3.3.1 Theoretical Formulation of TextMatch: Stabilizing Text Matching by Synthesis
$\widehat{\mathrm{PCI}}(a,b)=y_{b}-\sum_{c_{i}\in\bm{C}}w_{i}y_{i}=y_{b}-\sum_{c_{i}\in\bm{C}}\frac{m_{i}}{\sum_{c_{i}\in\bm{C}}m_{i}}y_{i}~{},$ (6)
@@ 3.3.3 Key Challenges and Mitigation Methods
We address several technical challenges below.
@@
Excluding the Mediators from Confounders Another key challenge to highlight is that the text variable we use for the confounder might accidentally include some mediator information. For example, the quality or performance of a paper could be expressed in the abstract, such as “we achieved 90% accuracy.” Therefore, we conduct a specific preprocessing procedure before feeding the text variable to the similarity function. For the $\mathrm{RemoveMediator}$ function in 19, we exclude all numerical expressions such as percentage numbers, as well as descriptions such as “state-of-the-art.” For generalizability, the essence of this step is a entanglement action to separate the confounder variable (in this case, the research content) and all the descendants of the treatment variable (in this case, mentions of the performance). For more complicated cases in future work, we recommend a separate disentanglement model to be applied here.
@@
Numerical Estimation Given the large number of papers, it is numerically challenging to aggregate the TCI from individual PCIs, because the number of follow-up papers for a study can be up to tens of thousands, such as the 57,200 citations by 2023 for the ImageNet paper (Deng et al., 2009). To avoid extensively running PCI for all follow-up papers, we propose a new numerical estimation method using a carefully designed random paper subset.
@@ 4.2 Author-Identified Paper Impact
Table 1 reports the accuracy of our CausalCite metric, together with two existing citation metrics: citations, and SSHI citations (Valenzuela-Escarcega et al., 2015). See the detailed derivation of the accuracy scores in Section C.2. From this table, we can see that our CausalCite metric achieves the highest accuracy, 80.29%, which is 5 points higher than SSHI, and 9 points higher than the traditional citations.
@@ 4.3 Test-of-Time Paper Analysis
In Table 2, we show the correlations of various metrics with the ToT awards. In this table, CausalCite achieves the highest correlation of 0.639, which is +30.14% better than that of citations. Furthermore, we visualize the correspondence of our metric and ToT, and observe a substantial difference between the CausalCite distributions of ToT vs. non-ToT papers in Figure 4. We also show three examples of ToT papers in Figure 4, where the ToT papers differ from the non-ToT papers by one or two orders of magnitude.
@@ 5.1 Do Best Papers Have High Causal Impact?
5.2 What Is the Nature of the CausalCite Distribution?
@@ 6 Related Work
With the increasing availability of large datasets and the advent of digital technologies, new opportunities for bibliometric analysis have emerged. Iqbal et al. (2021) highlighted the role of NLP and machine learning in enhancing in-text citation analysis. Similarly, Umer et al. (2021) explored the use of textual features and SMOTE resampling techniques in scientific paper citation analysis. Jebari et al. (2021) analyzed citation context to detect research topic evolution, showcasing data analysis for scientific discourse. Chang et al. (2023) explored augmenting citations in scientific papers with historical context, offering a novel perspective on citation analysis. Manghi et al. (2021) introduced scientific knowledge graphs, an innovative method for evaluating research impact. Bittmann et al. (2021) explored statistical matching in bibliometrics, discussing its utility and challenges in post-matching analysis. The use of AI in bibliometric analysis is highlighted in research by Chubb et al. (2022) and the systematic review of AI in information systems by Collins et al. (2021). Network analysis approaches, as discussed by Chakraborty et al. (2020) in the context of patent citations and by Dawson et al. (2014) in learning analytics, further illustrate the diverse applications of advanced methodologies in understanding citation patterns.
@@ 7 Conclusion
There are several limitations for our work. For example, as mentioned previously, our metric has a high computational budget. Future work can explore more efficient optimization methods. Also, we model the content of the paper by its title and abstract, it could also be possible for future work to benefit from modeling the full text, given appropriate license permissions.
@@
Data Collection and Privacy The data used in this work are all from Open Source Semantic Scholar data, with no user privacy concerns. The potential use of this work is for finding papers that are unique and innovative but do not get enough citations due to loack of popularity or awareness of the field. This metric can act as an aid when deciding impact of papers, but we do not suggest its usage without expert involvement. Through this work, we are not trying to demean or criticize anyone’s work we only intend to find more papers that have made a valuable contribution to the field.
CS-Centric Perspective The authors of this paper work in Computer Science (mostly Machine Learning) hence a lot of analysis done on the quality of papers that required sanity checks are done on ML papers. The conferences selected for doing the ToT evaluation were also CS Top conferences, hence they might have induced some biases. The metric in general has been created generically and should be applicable to other domains as well, the Author Identified Most Influential Papers study is also done on a generalized dataset, but we encourage readers in other disciplines to try out the metric on papers from their field.
@@
This project originates as part of the AI Scholar series of projects that Zhijing Jin started since 2021, as she identified that causal inference over papers is a valuable research setting with sufficient data and rich causal phenomena. Bernhard Schölkopf came up with the formulation that the action of citation itself has a causal nature, and can thus be formulated as a causal inference question. Zhijing, Bernhard, and Siyuan Guo settled down the overall project design.
@@ A.3 Experiment to Select the Best Embedding Method
When selecting the text encoder for our TextMatch method, we compare among the three LLMs pre-trained on scientific papers, SciBERT, MPNet, and SPECTER. Specifically, we conduct a small-scale experiment to see how much the similarities scores based on the embedding of each model align with human annotations. As for the annotation process, we first collect a set of random papers, and for each such paper (which we call a pivot paper), we identify ten papers, from the most similar to the least, with monotonically decreasing similarity. We collect a total of 100 papers consisting of ten such collections, for which we show an example in Table 6. Then we see how the resulting similarity scores conform to this order by deducting the percentage of papers that are out of place in the ranking.
@@ C.2 Additional Information for the Author-Identified Paper Impact Experiment
- A Quantitative Analysis and Performance Study for Similarity-Search Methods in High-Dimensional Spaces 0 3.519 1777 156
- Optimizing bitmap indices with efficient compression 0 3.519 375 40
- Data Warehouses And Olap: Concepts, Architectures And Solutions 0 3.526 187 11
- Histogram-aware sorting for enhanced word-aligned compression in bitmap indexes 0 3.543 17 1
- CubiST++: Evaluating Ad-Hoc CUBE Queries Using Statistics Trees 0 3.543 5 1
- Improving Performance of Sparse Matrix-Vector Multiplication 0 3.543 114 11
- Binary Gray Codes with Long Bit Runs 0 3.543 53 4
- Analysis of Basic Data Reordering Techniques 0 3.543 16 1
- Tree Based Indexes Versus Bitmap Indexes: A Performance Study 0 3.543 24 0
- Secondary indexing in one dimension: beyond b-trees and bitmap indexes 0 3.543 10 1
- A comparison of five probabilistic view-size estimation techniques in OLAP 0 3.543 24 1
- Compression techniques for fast external sorting 0 3.543 16 0
- A Note on Graph Coloring Extensions and List-Colorings 0 3.543 33 1
- Using Multiset Discrimination to Solve Language Processing Problems Without Hashing 0 3.543 52 2
- Monotone Gray Codes and the Middle Levels Problem 0 3.543 80 5
- The Art in Computer Programming 0 3.543 9242 678
- An Efficient Multi-Component Indexing Embedded Bitmap Compression for Data Reorganization 0 3.543 8 2
- The LitOLAP Project: Data Warehousing with Literature 0 3.543 8 0
- Multi-resolution bitmap indexes for scientific data 0 3.583 96 3
- Notes on design and implementation of compressed bit vectors 0 3.583 81 12
- Compressing Large Boolean Matrices using Reordering Techniques 0 3.595 88 7
- Compressing bitmap indices by data reorganization 1 3.595 53 4
- Model 204 Architecture and Performance 0 3.635 238 10
- On the performance of bitmap indices for high cardinality attributes 1 3.654 196 10
- A performance comparison of bitmap indexes 0 3.655 86 9
- Minimizing I/O Costs of Multi-Dimensional Queries with Bitmap Indices 0 3.692 16 0
- Evaluation Strategies for Bitmap Indices with Binning 0 3.692 69 3
- C-Store: A Column-oriented DBMS 0 3.710 1241 111
- Byte-aligned bitmap compression 0 3.793 209 48
- Bit Transposed Files 0 3.837 84 10
- Space efficient bitmap indexing 0 4.011 96 16